OpenStudy (anonymous):

Emily has 64 books in her book collection. She has 20 mystery book. The rest is fiction. What is the fraction( in simplest form) for the fiction books.

OpenStudy (anonymous):

(64-20)/64

OpenStudy (anonymous):

reduce that fraction

OpenStudy (anonymous):

but aren't mystery books fiction?

OpenStudy (anonymous):

idk.

OpenStudy (anonymous):

44/64, 22/32, 11/16 is the answer

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

Divide by 2, or divide by 4 for a faster answer

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

mystery books dont have to b fiction!

OpenStudy (anonymous):

Haha, so very true

OpenStudy (anonymous):

its just easier to write 1 when they are :P

OpenStudy (anonymous):

44/64 =22/32 =11/16